osesin and 3 Given sin e 5 -7 37 sin B= 25 2 Find tan(20) <B< 27.
Answer #1

no au 2 As os. okt trigonon onomersie functions of are positive. Now sin(0) = 3 implies, cos e = Insino 11- 975 = 1/2 so, tom​​​​​​.

.

So we have,

tan(2θ) = 24/7